OpenStudy (liv1234):

What is the inverse of the given relation? y=3x+12 My answer is (x-12)/3

OpenStudy (liv1234):

@priyar

OpenStudy (priyar):

yep! correct!

OpenStudy (igreen):

y = 3x + 12 x = 3y + 12 x - 12 = 3y (x - 12)/3 = y Yep!

OpenStudy (mathwizzard3):

I believe so. I'm a little rusty on this though

OpenStudy (liv1234):

Thank you all! I really appreciate your help!

OpenStudy (priyar):

f^-1 (x) =(x-12)/3 would be more appropriate!

OpenStudy (priyar):

if y=f(x) that is!